### v2

@@ Abstract.
The challenge of representing evolutionary relationships in phylogenetic networks has been described as one of the key problems in the use of mathematics in biology [8], and the last decade has seen significant investment in exploring their modelling, inference, and mathematics. The motivation is clear: well-studied evolutionary processes, such as horizontal gene transfer, endosymbiosis, and hybridization, are not able to be represented by a phylogenetic tree [11, 12, 10].
The difficulty is that, in contrast to phylogenetic trees, there are infinitely many phylogenetic networks that relate a given set of taxa appearing at the leaves of the network. This makes inference decidedly harder: it is not even theoretically possible to search the whole space for optimal solutions.
@@
Many families of phylogenetic network have been defined, each seeking to capture either biological intuition or desirable mathematical features. Normal networks are one such class, and are a subclass of the tree-child networks. A network is tree-child if every internal vertex in the network has a child that is a tree vertex [7]. Each vertex $v$ in a tree-child network has the property that there is a leaf $x$ for which every path from the root to $x$ passes through $v$ [7, Lemma 2]. This property is called “visibility” because it implies that the evolutionary signal from $v$ is visible in a leaf, $x$ (see Figure 2 for an example).
@@
The class of tree-child networks is in turn contained in the classes of orchard networks [13] and tree-sibling networks [6], which are briefly defined here. The orchard networks are those that can be progressively reduced via “cherry reductions” and “reticulated cherry reductions” at the leaves, to a single vertex. They have the attractive property of being characterised by their “ancestral profiles”, which are a set of tuples for the leaves of the network that enumerate the number of paths from each vertex to each leaf. Tree-sibling networks are those for which every reticulation vertex has a tree vertex as a sibling (a vertex that shares the same parent). Both the orchard and tree-sibling classes in turn are contained in the class of tree-based networks, which are networks that have a spanning tree whose leaves are those of the networks [17], and the class of labellable networks, which are those whose internal vertices can be deterministically labelled [18] (many classes, including normal networks, can be characterised in the labellable framework [15]).
@@
Figure 3. Inclusion relationships among some classes of rooted binary phylogenetic networks, where edges denote inclusion of a class in the class above. All class names are classes of networks, except for Trees and Caterpillar trees shown in dashed boxes at the bottom. This figure is adapted from [25, Fig 12], showing a subset of networks from that figure while including some additional classes and newly known relationships. The additional classes shown are the recently-defined labellable and spinal network classes [18, 15], and the hybridization class. To connect spinal networks with trees, Caterpillar trees have been added below, as a subclass of Trees. The new relationships are that hybridization networks are all normal [30, Prop 10.12], that the HGT and orchard network classes are identical [32], and that fold-unfold (FU)-stable networks are labellable (this follows from [23, Thm 1] and the characterization of labellable networks [18, Thm 3.3]). Note that stack-free and reticulation-visible networks are not labellable, because there are many such networks that violate [18, Thm 3.3]. For definitions and references for the other relationships in the figure, see [25].
The key sources of non-tree-like evolutionary processes in evolutionary histories (giving reticulations in the phylogenetic networks) are hybridization, and horizontal (or lateral) gene transfer (HGT). Broadly speaking, the former are more common in eukaryotes (plants, animals, fungi), the latter in prokaryotes (bacteria, archaea). Networks with these specific processes in mind can be defined as trees with additional types of event formalised, and these constraints provide more specific restrictions that help narrow the focus within this ecosystem of classes (see Fig. 4 for examples).
For instance, it turns out that binary hybridization networks are normal networks [30, Prop. 10.12]. If one is seeking a network that describes the evolutionary history of a set of eukaryotic species, this is the sort of network that one is most likely to want to reconstruct. If on the other hand, one is seeking a network describing prokaryotic evolution, then a horizontal gene transfer (HGT) network might be more relevant. HGT networks allow more (mathematical) freedom, because each event only requires one arc. Despite this, they have recently been shown to be precisely the orchard networks [32], which contains the class of normal networks.
@@
At a larger scale than substructures on three or four leaves, the trees displayed by a network (which have the same leaf-set as the network), along with their multiplicities, are important, because they are often interpreted as representing the evolutionary histories of individual genes within the population. For normal networks, early results showed that in the binary case (in which each internal vertex has degree 3), that they are among those that can be reconstructed from their displayed trees [36]. But there are stronger facts about displayed trees for normal networks. For instance, normal networks display any given tree at most once [35, 9]. This means that for binary normal networks with $r$ reticulations, there are exactly $2^{r}$ trees displayed in the network. Most powerfully, a normal phylogenetic network is uniquely determined by these $2^{r}$ trees that it displays, and indeed can be reconstructued from those trees by a polynomial time algorithm [4].
@@
Figure 5. A phylogenetic network $N$, from Figure 1, and its normalisation. Vertices are shown solid if they are visible. The non-visible vertices in $N$ are those vertices $v$ with the property that for any leaf $x_{i}$ there is a path from the root of $N$ to $x_{i}$ that does not pass through $v$. Thus, tracing back from any extant leaf to the root, it is possible to never “see” a non-visible vertex as there is always a path that avoids it. The normalisation of $N$ on the right shows all visible vertices in $N$ except $v_{9}$, which is suppressed since it would be degree 2 in the cover digraph of the visible vertices of $N$.
@@ Conclusion
Phylogenetics is the science of inferring ancestral relationships among a set of present-day species. With such a project, and in the presence of reticulate events such as hybridization and horizontal gene transfer, there can inevitably be information whose signal is irretrievably lost. As the mathematical properties of phylogenetic networks continue to be developed, we can expect clarity to emerge about which classes of network are recoverable from different kinds of data, and we can expect the development of robust methods to make inferences. The current evidence, as summarized here, seems to be indicating that the class of normal networks may provide a framework for this emergent clarity, and for inference methods that respect the information constraints imposed by reticulation. The further development of methods for the reconstruction of normal networks, and results that develop their mathematical properties, seems a most promising direction for future research.

### v3

@@ Abstract.
The need to be able to represent evolutionary relationships that cannot be shown by familiar phylogenetic trees has been recognised for half a century now [35, 24, 13]. These relationships, such as horizontal gene transfer (HGT), endosymbiosis, and hybridization, play a significant evolutionary role, and so how to infer and represent them, using phylogenetic networks, is increasingly important [11, 12, 10]. How exactly to do that has been described as one of the key problems in the use of mathematics in biology [8].
The last couple of decades has seen significant investment in exploring the modelling, inference, and mathematics of phylogenetic networks. The difficulty is that, in contrast to phylogenetic trees, there are infinitely many phylogenetic networks that relate a given set of taxa appearing at the leaves of the network. This makes inference decidedly harder: it is not even theoretically possible to search the whole space for optimal solutions.
@@
Many families of phylogenetic network have been defined to date, each seeking to capture either biological intuition or desirable mathematical features. “Normal” networks are one such class, and are a subclass of the “tree-child” networks. They were first defined, and given that name, by Willson in 2010 [41].
A network is tree-child if every internal vertex in the network has a child that is a tree vertex or leaf [7]. Each vertex $v$ in a tree-child network has the property that there is a leaf $x$ for which every path from the root to $x$ passes through $v$ [7, Lemma 2]. This property is called “visibility” because it implies that the evolutionary signal from $v$ is visible in a leaf, $x$ (see Figure 2 for an example).
@@
The class of tree-child networks is in turn contained in the classes of orchard networks [14] and tree-sibling networks [6], which are briefly defined here. The orchard networks are those that can be progressively reduced via “cherry reductions” and “reticulated cherry reductions” at the leaves, to a single vertex. They have the attractive property of being characterised by their “ancestral profiles”, which are a set of tuples for the leaves of the network that enumerate the number of paths from each vertex to each leaf [14]. Tree-sibling networks are those for which every reticulation vertex has a tree vertex as a sibling (a vertex that shares the same parent). Both the orchard and tree-sibling classes in turn are contained in the class of tree-based networks, which are networks that have a spanning tree whose leaves are those of the networks [18], and the class of labellable networks, which are those whose internal vertices can be deterministically labelled [19] (many classes, including normal networks, can be characterised in the labellable framework [16]).
Networks that model only specific processes, such as HGT or hybridization, can be formally defined as trees with additional types of event permitted, and these constraints provide more specific restrictions that help narrow the focus within this ecosystem of classes. HGT networks and hybridization networks, represent the named event as an almost instantaneous (horizontal) effect on the network, as shown in Fig. 3 (for example, see [21]). In contrast, other families of network usually represent reticulate events with ambiguous timing, which allows uncertainty in the representation (that is, reticulate arrows need not be horizontal), as shown in the example in Fig. 1.
@@
Figure 4. Inclusion relationships among some classes of rooted binary phylogenetic networks, where edges denote inclusion of a class in the class above. All class names are classes of networks, including the special classes of Trees and Caterpillar trees shown in dashed boxes at the bottom. This figure is adapted from [29, Fig 12], showing a subset of networks from that figure while including some additional classes and newly known relationships. The additional classes shown are the recently-defined labellable and spinal network classes [19, 16], and the hybridization class. To connect spinal networks with trees, Caterpillar trees have been added below, as a subclass of Trees. The new relationships are that hybridization networks are all normal [36, Prop 10.12], that the HGT and orchard network classes are identical [39], and that fold-unfold (FU)-stable networks are labellable (this follows from [27, Thm 1] and the characterization of labellable networks [19, Thm 3.3]). Note that stack-free and reticulation-visible networks are not labellable, because there are many such networks that violate [19, Thm 3.3]. For definitions and references for the other relationships in the figure, see [29].
For instance, it turns out that binary hybridization networks are normal networks [36, Prop. 10.12]. If one is seeking a network that describes the evolutionary history of a set of eukaryotic species, this is the sort of network that one is most likely to want to reconstruct. If on the other hand, one is seeking a network describing prokaryotic evolution, or even tumor evolution (see e.g [38]), then a horizontal gene transfer (HGT) network might be more relevant. HGT networks allow more (mathematical) freedom, because each event only requires one arc. Despite this, they have recently been shown to be precisely the orchard networks [39], which contains the class of normal networks.
@@
At a larger scale than substructures on three or four leaves, the trees displayed by a network (which have the same leaf-set as the network), along with their multiplicities, are important, because they are often interpreted as representing the evolutionary histories of individual genes within the population. For normal networks, early results showed that in the binary case (in which each internal vertex has degree 3), they are among those that can be reconstructed from their displayed trees [43]. But there are stronger facts about displayed trees for normal networks. For instance, normal networks display any given tree at most once [42, 9]. This means that for binary normal networks with $r$ reticulations, there are exactly $2^{r}$ trees displayed in the network. Most powerfully, a normal phylogenetic network is uniquely determined by these $2^{r}$ trees that it displays, and indeed can be reconstructued from those trees by a polynomial time algorithm [4].
@@
Figure 5. A phylogenetic network NN, from Figure 1, and its normalisation. Vertices are shown solid if they are visible. The non-visible vertices in NN are those vertices $v$ with the property that for any leaf $x_{i}$ there is a path from the root of NN to $x_{i}$ that does not pass through $v$. Thus, tracing back from any extant leaf to the root, it is possible to never “see” a non-visible vertex as there is always a path that avoids it. The normalisation of NN on the right shows all visible vertices in NN except $v_{9}$, which is suppressed since it would be degree 2 in the cover digraph of the visible vertices of NN.
@@ Conclusion
Phylogenetics is the science of inferring ancestral relationships among a set of present-day species. With such a project, and in the presence of reticulate events such as hybridization and horizontal gene transfer, there can inevitably be information whose signal is irretrievably lost. As the mathematical properties of phylogenetic networks continue to be developed, we can expect clarity to emerge about which classes of network are recoverable from different kinds of data, and we can expect the development of robust methods to make inferences. The current evidence, as summarized here, seems to be indicating that the class of normal networks may provide a framework for this emergent clarity, and for inference methods that respect the information constraints imposed by reticulation.
There will certainly be more nuance to come. We have not discussed details of several dimensions of network inference, making at most a passing reference to implicit networks, or the related semidirected networks [23, 31], and have not gone far into networks that arise from different types of data or processes (aside from a brief mention of HGT and hybridization networks). The further development of methods for the reconstruction of normal networks, and results that develop their mathematical properties, seems a most promising direction for future research.